Can someone please help me on this one, I have no clue of what to do! Thank you
Consider a die with eight sides, marked one, two, three, and so on. Assuming equally likely outcomes, find the probability that the sum of two dice is the given number. (Enter the answer either as a fracion or as a decimal rounded to three places.
(a) P(3)=
(b) P(4)=
(c) P(7)=
(d) P(10)=
(e) P(12)=
(f) P(13)=

Answer by stanbon(75887)

Consider a die with eight sides, marked one, two, three, and so on. Assuming equally likely outcomes, find the probability that the sum of two dice is the given number. (Enter the answer either as a fracion or as a decimal rounded to three places.
---------
This is "what to do".
Sketch an 8 by 8 rectangle with rows labled 1-8
and columns labled 1-8
---
Enter the sum of each row-label + column-label in
the row/column position:
You should get the following in the 1st few rows:
2/3/4/5/6/7/8/9/10
3/4/5/6/7/8/9/10/11
4/5/6/7/8/9/10/11/12
etc.
-----------
(a) P(3)=
You see there are a total of 8*8 = 64 sums
and 2 of the sums is "3".
So P(3) = 2/64
----
(b) P(4)= 3/64
etc.
----
(c) P(7)=
(d) P(10)=
(e) P(12)=
(f) P(13)=
=================
Cheers,
Stan H.